Candice ask 100 students put Chris plate over there like best out of ready salted change an arsehole tomato footage of the students are boys 7 to use a boys like cheese and onion the best 15 girls like to resold the best 23 out of 39 students you like salt and vinegar best for girlsCandice ask 100 students put Chris plate over there like best out of ready salted change an arsehole tomato footage of the students are boys 7 to use a boys like cheese and onion the best 15 girls like to resold the best 23 out of 39 students you like salt and vinegar best of girls work out the number of students who like to
1
Expert's answer
2019-04-09T11:17:39-0400

7/100*100% = 7% of the best salted tomato made by boys. 15/100*100%=15 % of the number of students-girls resale your tomatoes. 23/100*100% = 23 % of the students are students who have used vinegar and salt